# To Buy an Electric Vehicle or Not? A Bayesian Analysis of Consumer Intent in the United States

### v1

@@ 4.1 Results from Ordinal Probit Models
Confidence in charging infrastructure, assessed in the 2023 survey, is one of the strongest predictors of EV interest. Respondents who are very or extremely confident that the US will build the necessary charging infrastructure are 19.1 percentage points more likely to express strong interest in purchasing EVs and 18.8 percentage points less likely to report no interest compared to those who report lower confidence. This finding aligns with prior research, such as Carley et al. (2013) and Li et al. (2017). As shown in Table 1, 82% of respondents lack confidence that US will develop the infrastructure needed to support EV adoption. Taken together, these results suggest that skepticism about charging infrastructure is likely to be a significant barrier to broader adoption, highlighting the importance of both improving infrastructure and clearly communicating its availability, reliability, and expansion plans to the public.
@@
Table 3: Covariate effects (CE) from ordinal probit models. CE for variables not statistically different from zero are marked with $\times$.
@@
$\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4) $\Delta$P(y=1) $\Delta$P(y=2) $\Delta$P(y=3) $\Delta$P(y=4)
EV Info $-0.0851$ $-0.0388$ $0.0217$ $0.1022$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
Env Better $-0.1955$ $-0.0565$ $0.0970$ $0.1551$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
EV Infra ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . $-0.1880$ $-0.0553$ $0.0526$ $0.1908$
EV Owner $-0.1321$ $-0.0849$ $0.0051$ $0.2119$ $-0.1749$ $-0.0794$ $0.0231$ $0.2312$ $-0.1814$ $-0.0513$ $0.0421$ $0.1905$
GACC (TM) $0.1971$ $0.0453$ $-0.0945$ $-0.1478$ $0.3341$ $0.0133$ $-0.1507$ $-0.1968$ $0.3170$ $-0.0153$ $-0.1395$ $-0.1622$
GACC (RA) $0.1180$ $0.0391$ $-0.0490$ $-0.1082$ $0.1649$ $0.0267$ $-0.0639$ $-0.1278$ $0.1282$ $0.0079$ $-0.0486$ $-0.0875$
Age $<50$ $-0.0326$ $-0.0130$ $0.0099$ $0.0356$ $-0.0410$ $-0.0090$ $0.0129$ $0.0371$ $-0.0371$ $-0.0036$ $0.0126$ $0.0282$
Female $\times$ $\times$ $\times$ $\times$ $0.0268$ $0.0055$ $-0.0082$ $-0.0240$ $0.0301$ $0.0027$ $-0.0100$ $-0.0228$
Grad & Above $-0.0815$ $-0.0322$ $0.0276$ $0.0860$ $-0.1139$ $-0.0240$ $0.0378$ $0.1001$ $-0.1264$ $-0.0130$ $0.0445$ $0.0949$
Some College $-0.0373$ $-0.0152$ $0.0099$ $0.0426$ $-0.0541$ $-0.0118$ $0.0156$ $0.0503$ $-0.0517$ $-0.0052$ $-0.0166$ $0.0403$
Married ..absent... . ..absent... . ..absent... . ..absent... . $-0.0154$ $-0.0029$ $0.0050$ $0.0134$$\times$ $\times$ $\times$ $\times$
Metropolitan $-0.0534$ $-0.0179$ $0.0191$ $0.0522$ $-0.0662$ $-0.0099$ $0.0232$ $0.0528$ $-0.0512$ $-0.0027$ $0.0183$ $0.0356$
Midwest $0.0162$ $0.0061$ $-0.0051$ $-0.0172$$\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
West $-0.0378$ $-0.0164$ $0.0101$ $0.0441$ $-0.0457$ $-0.0109$ $0.0135$ $0.0431$ $-0.0454$ $-0.0053$ $0.0148$ $0.0360$
US Born $0.0416$ $0.0186$ $-0.0106$ $-0.0497$ $0.0777$ $0.0217$ $-0.0214$ $-0.0781$ $0.0551$ $0.0071$ $-0.0178$ $-0.0444$
Black $0.0614$ $0.0200$ $-0.0220$ $-0.0594$ $0.0349$ $0.0058$ $-0.0116$ $-0.0291$ $0.0460$ $0.0026$ $-0.0160$ $-0.0326$
@@
Democrat $-0.1211$ $-0.0578$ $0.0403$ $0.1386$ $-0.1569$ $-0.0497$ $0.0513$ $0.1553$ $-0.1882$ $-0.0353$ $0.0677$ $0.1558$
Ind & Others $-0.0808$ $-0.0310$ $0.0219$ $0.0898$ $-0.0890$ $-0.0187$ $0.0266$ $0.0811$ $-0.1091$ $-0.0093$ $0.0356$ $0.0827$
$\Delta$I=\$$=\$50,000$$\times$ $\times$ $\times$ $\times$ $-0.0223$ $-0.0051$ $0.0066$ $0.0209$ $-0.0363$ $-0.0045$ $0.0116$ $0.0291$
Across all three surveys, previous EV ownership is consistently linked with a higher interest in purchasing an EV. Respondents who have owned an EV or hybrid vehicle are 19.0–23.1 percentage points more likely to report strong interest in buying an EV and 13.2–18.1 percentage points less likely to express no interest compared to those without prior ownership. Other studies such as Carley et al. (2013) and Buhmann and Criado (2023) have also documented this pattern, suggesting that firsthand experience reduces uncertainty and fosters positive perceptions of EVs. However, selection effects may also be at play, as these individuals are likely early adopters with a strong pre-existing interest in the technology, making them more engaged with the EV market.
Respondents who think the government is doing too much to address climate change (GACC, TM) are 14.8–19.7 percentage points less likely to express strong interest in purchasing EVs and 19.7–33.4 percentage points more likely to report no interest. Those who believe the government is doing the right amount (GACC, RA) also express lower interest, though the effects are smaller.
Respondents under age 50 are 2.8–3.7 percentage points more likely to express strong interest and 3.3–4.1 percentage points less likely to report no interest in purchasing EVs compared to older respondents. Our finding aligns with Zhao et al. (2022), Carley et al. (2013), Krause et al. (2013), Scherrer (2023), and Mamkhezri (2025). For instance, Zhao et al. (2022) finds that EV buyers in Shanghai are predominantly younger, with most between the ages of 31 and 40.
Results from the 2022 and 2023 surveys also show that female respondents are modestly less likely to express strong interest and more likely to report no interest than males. The covariate effects on strong interest decrease between 2.3–2.4 percentage points; while the covariate effect for no interest increase between 2.7–3.0 percentage points. This finding is consistent several studies, including Carley et al. (2013), Krause et al. (2013), Scherrer (2023), Buhmann and Criado (2023), and Mamkhezri (2025), which also find that men in their samples exhibit a stronger preference for EVs.
Higher education is also linked to greater EV interest. Respondents with a graduate degree are 8.6–10.0 percentage points more likely to report strong interest and 8.1–12.6 percentage points less likely to express no interest compared to those with a high school diploma or less. Similar, though smaller, effects are observed for respondents with some college education. The positive relationship between education and EV interest is similarly documented in Carley et al. (2013), Krause et al. (2013), Buhmann and Criado (2023), and Mamkhezri (2025).
Metropolitan residents are 3.6–5.3 percentage points more likely to report strong interest and 5.1–6.6 percentage points less likely to express no interest. This finding aligns with Buhmann and Criado (2023) and Mamkhezri (2025), who also report EV preference to be higher among urban residents. Moreover, respondents in the West are 3.6–4.4 percentage points more likely to report strong interest compared to those in the Northeast, though those in the Midwest and South show no significant difference from their Northeastern counterparts. This pattern likely reflects California’s leadership in EV adoption, driven in part by policies such as ZEV mandates and the Clean Vehicle Rebate Program, along with a robust charging network. [^5] See https://www.eia.gov/todayinenergy/detail.php?id=61082
US-born respondents are 4.4–7.8 percentage points less likely to express strong interest and 4.2–7.8 percentage points more likely to express no interest compared to non-US-born respondents, possibly reflecting differences in cultural attitudes, driving patterns, or perceptions about the cost of vehicle ownership. Such a difference in perception about EVs is also evident in the racial background of respondents. Black respondents are 2.9–5.9 percentage points less likely to express strong interest and 3.5–6.1 percentage points more likely to express no interest, as compared to White respondents, likely reflecting structural barriers, such as limited access to charging infrastructure (Hsu and Fingerman, 2021). Other non-white respondents show similar preferences to White respondents.
Political affiliation is a consistent and significant predictor of EV purchase intent. Democrats are 13.9–18.8 percentage points more likely to report strong interest in EVs and 12.1–18.8 percentage points less likely to report no interest compared to Republicans. Similarly, Independents and others are also about 8 percentage points more likely to express strong interest in purchasing EVs relative to Republicans. These findings align with Mamkhezri (2025) and likely reflect Democrats’ stronger perceptions of EVs as environmentally responsible compared to Republicans, as documented by Sintov et al. (2020). They also reflect the starkly contrasting policy approaches of the two major political parties, particularly during the tenures of Presidents Biden and Trump. During his first term (2017–2021), President Trump’s administration rolled back Obama-era fuel efficiency standards, slowing regulatory pressure on automakers to transition toward EVs. In contrast, the Biden administration adopted an aggressive pro-EV stance as part of its broader climate agenda, prioritizing stricter fuel efficiency and emissions standards and allocating \$7.5 billion through the Infrastructure Investment and Jobs Act to develop a nationwide network of 500,000 EV charging stations.
@@
Conditional on other factors, an additional \$50,000 in income is either not significantly associated with EV interest or changes the probability of strong interest (no interest) by a meagre 2.1–2.9 (2.2–3.6) percentage points. Our results find support in Zhao et al. (2022) and Mamkhezri (2025), who report greater support for EVs among higher-income individuals. The small covariate effect for income may reflect the growing availability of EVs across various market segments and the presence of subsidies for low-income consumers, which reduce affordability barriers compared to earlier market years.
@@
Overall, the results show that interest in EVs varies based on environmental attitudes, experience with EVs, confidence in charging infrastructure, and demographic factors. Some patterns likely reflect personal preferences, while others may reflect barriers such as limited awareness or charging access.
@@ 4.2 Results from Ordinal Quantile Models
We now present the results from ordinal quantile models (Rahman, 2016) that shift attention from average probabilities of outcomes to the probabilities of outcomes at specific quantiles, conditional on the covariates. In particular, Table 4 presents the posterior means and standard deviations of the parameters from the ordinal quantile models along with the logarithm of marginal likelihood, estimated separately for each survey data. Table 5 reports the corresponding covariate effects. The covariate effects at the 50th percentile align closely with those from the ordinal probit model. However, across all surveys, the covariate effects at the 20th and 80th quantiles are noticeably different compared to those from the ordinal probit model. As such, we find compelling evidence that the impact of covariates on the intent to purchase EV is not uniform, but varies across the distribution of EV purchase intent. This highlights the value of quantile modeling for capturing these variations. Below, we emphasize the key differences across quantiles.
While having heard or read extensively about EVs corresponds to higher purchase intent, the relationship is stronger among respondents more inclined toward adoption. In the 2021 survey and at the 80th percentile of EV purchase intent, respondents who have heard or read extensively about EVs are 12.6 percentage points more likely to express strong intent to purchase compared to those who have read or heard little to nothing. At the 50th and 20th percentiles of EV purchase intent, the difference is 9.9 and 6.1 percentage points, respectively. The smaller covariate effect of prior exposure to EV information at the 20th percentile, suggests that simply raising general awareness may not be enough to persuade consumers with low purchase interest. In contrast, for those already inclined toward adoption, providing practical, consumer-focused information—such as total cost of ownership, available incentives, and charging logistics—may be effective in influencing purchase decisions.
Similarly, the relationship between confidence that the US will build the necessary charging infrastructure (measured in the 2023 survey) and EV purchase intent varies across the distribution of EV purchase intent. At the 80th percentile, respondents who are confident in this development are 20.0 percentage points more likely to express strong intent to purchase compared to those who are not confident. The corresponding differences are 18.1 percentage points at the median and 13.3 percentage points at the 20th percentile. Conversely, at the 80th percentile, those confident in infrastructure development (relative to those who are not) are 14.5 percentage points less likely to express no interest, with differences of 18.3 and 18.6 percentage points at the median and 20th percentile, respectively. These findings suggest that confidence in infrastructure development strengthens commitment among those already inclined toward adoption and reduces outright rejection among those with lower purchase intent. Enhancing the affordability, accessibility, and reliability of charging infrastructure—and effectively communicating these efforts—can build consumer confidence and promote broader EV adoption.
Table 4: Results from quantile ordinal model: posterior mean (Mean) and standard deviation (Std) of the parameters, along with the logarithm of marginal likelihood, from April 2021, May 2022, and May–Jun 2023 survey data.
@@
Intercept $-3.60$ $0.25$ $-0.58$ $0.17$ $2.60$ $0.25$ $-2.15$ $0.30$ $0.80$ $0.21$ $4.64$ $0.29$ $-2.81$ $0.30$ $-0.05$ $0.20$ $2.77$ $0.28$
EV Info $1.28$ $0.08$ $0.98$ $0.06$ $1.64$ $0.08$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
Env Better $2.52$ $0.11$ $1.80$ $0.07$ $2.85$ $0.10$.absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... .
EV Infra ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . $2.83$ $0.13$ $1.92$ $0.10$ $2.71$ $0.14$
@@
GACC (TM) $-2.81$ $0.13$ $-1.96$ $0.09$ $-2.49$ $0.13$ $-3.79$ $0.16$ $-2.63$ $0.10$ $-3.70$ $0.14$ $-4.00$ $0.17$ $-2.41$ $0.10$ $-3.45$ $0.14$
GACC (RA) $-1.79$ $0.10$ $-1.24$ $0.07$ $-1.55$ $0.11$ $-1.94$ $0.12$ $-1.48$ $0.08$ $-2.03$ $0.12$ $-1.77$ $0.12$ $-1.15$ $0.08$ $-1.48$ $0.12$
Age$<50$ $0.40$ $0.08$ $0.36$ $0.05$ $0.58$ $0.08$ $0.50$ $0.09$ $0.37$ $0.07$ $0.63$ $0.10$ $0.42$ $0.11$ $0.34$ $0.07$ $0.55$ $0.10$
Female $-0.11$ $0.08$ $-0.02$ $0.05$ $-0.03$ $0.07$ $-0.36$ $0.10$ $-0.22$ $0.06$ $-0.33$ $0.09$ $-0.49$ $0.10$ $-0.31$ $0.07$ $-0.27$ $0.09$
Grad & Above $1.42$ $0.14$ $-0.93$ $0.08$ $1.17$ $0.13$ $1.93$ $0.16$ $1.04$ $0.10$ $1.19$ $0.13$ $1.94$ $0.16$ $1.09$ $0.10$ $1.43$ $0.13$
@@
Married ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . ..absent... . $0.22$ $0.13$ $0.12$ $0.08$ $0.19$ $0.12$ $0.02$ $0.13$ $0.08$ $0.08$ $0.12$ $0.11$
@@
US Born $-0.63$ $0.12$ $-0.49$ $0.09$ $-0.80$ $0.14$ $-0.89$ $0.15$ $-0.67$ $0.10$ $-1.25$ $0.16$ $-0.91$ $0.16$ $-0.45$ $0.11$ $-0.67$ $0.15$
Black $-1.23$ $0.15$ $-0.70$ $0.10$ $-0.56$ $0.14$ $-0.86$ $0.20$ $-0.25$ $0.14$ $-0.06$ $0.18$ $-0.94$ $0.16$ $-0.42$ $0.11$ $-0.36$ $0.16$
Other Races $0.08$ $0.10$ $0.14$ $0.07$ $0.25$ $0.11$ $-0.19$ $0.14$ $0.07$ $0.10$ $0.22$ $0.13$ $-0.30$ $0.15$ $0.00$ $0.10$ $0.02$ $0.14$
@@
$\ln$ ML $-12871$ $-12529$ $-12901$ $-8563$ $-8357$ $-8624$ $-8237$ $-7981$ $-8194$
Table 5: Covariate effects (CE) from quantile analysis. CE for variables not statistically different from zero are marked with $\times$.

### v2

@@ 4.1 Results from Ordinal Probit Models
Confidence in charging infrastructure, assessed in the 2023 survey, is one of the strongest predictors of EV interest. Respondents very or extremely confident that the US will build the necessary infrastructure are 19.1 (18.8) percentage points more likely to express strong (no) interest in purchasing EVs, compared to those who report lower confidence. This finding complements research, such as Carley et al. (2013) and Li et al. (2017), which highlight the importance of existing infrastructure.
Respondents who think the government is doing too much to address climate change (GACC, TM) are 14.8–19.7 percentage points less likely to express strong interest in purchasing EVs and 19.7–33.4 percentage points more likely to report no interest. Those who believe the government is doing the right amount (GACC, RA) also express lower interest, though to a lesser extent. This relationship has not, to our knowledge, been examined in prior work.
Across all surveys, previous EV ownership is consistently linked with a higher purchase interest. Respondents who have owned an EV or hybrid are 19.0–23.1 percentage points more likely to report strong interest in buying an EV and 13.2–18.1 percentage points less likely to express no interest than those without prior ownership. This pattern, also noted by Carley et al. (2013) and Buhmann and Criado (2023), suggests that firsthand experience reduces uncertainty and fosters positive perceptions of EVs. However, selection effects may also be at play, as these individuals are likely early adopters with a strong pre-existing interest in EVs.
Respondents under age 50 are 2.8–3.7 percentage points more likely to express strong interest and 3.3–4.1 percentage points less likely to report no interest in purchasing EVs than older respondents. Our finding aligns with Zhao et al. (2022), Carley et al. (2013), Krause et al. (2013), Scherrer (2023), and Mamkhezri (2025). For instance, Zhao et al. (2022) finds that EV buyers in Shanghai are mostly aged 31-40.
Results from the 2022 and 2023 surveys also show that female respondents are are 2.3–2.4 percentage points less likely to express strong interest, and 2.7–3.0 percentage points more likely to report no interest, than males. This finding is consistent with Carley et al. (2013), Krause et al. (2013), Scherrer (2023), Buhmann and Criado (2023), and Mamkhezri (2025), which also find that men show stronger preference for EVs.
Higher education is also linked to greater EV interest. Respondents with a graduate degree are 8.6–10.0 percentage points more likely to report strong interest and 8.1–12.6 percentage points less likely to express no interest than those with a high school diploma or less. Similar, though smaller, effects are observed for respondents with some college education. Such a relationship is similarly documented in Carley et al. (2013), Krause et al. (2013), Buhmann and Criado (2023), and Mamkhezri (2025).
Metropolitan residents are 3.6–5.3 percentage points more likely to report strong interest and 5.1–6.6 percentage points less likely to express no interest. This finding aligns with Buhmann and Criado (2023) and Mamkhezri (2025), who also report higher EV preference among urban residents. Moreover, respondents in the West are 3.6–4.4 percentage points more likely to report strong interest than those in the Northeast; those in the Midwest and South show no significant difference from Northeastern respondents. This pattern likely reflects California’s leadership in EV adoption, driven in part ZEV mandates, the Clean Vehicle Rebate Program, and a robust charging network. [^8] See https://www.eia.gov/todayinenergy/detail.php?id=61082
US-born respondents are 4.4–7.8 percentage points less likely to express strong interest and 4.2–7.8 percentage points more likely to express no interest than non-US-born respondents. Moreover, Black respondents are 2.9–5.9 percentage points less likely to express strong interest and 3.5–6.1 percentage points more likely to express no interest, as compared to White respondents. Other non-white respondents show similar preferences to White respondents. These differences may reflect cultural attitudes, driving patterns, cost perceptions, or structural barriers such as limited charging access (Hsu and Fingerman, 2021).
Political affiliation is a consistent and significant predictor of EV purchase intent. Democrats are 13.9–18.8 percentage points more likely to report strong interest in purchasing EVs and 12.1–18.8 percentage points less likely to report no interest than Republicans. Independents and others are about 8 percentage points more likely to express strong interest than Republicans. These findings align with Mamkhezri (2025) and likely reflect Democrats’ stronger perceptions of EVs as environmentally responsible, as documented by Sintov et al. (2020). They also reflect the stark policy differences between the two major political parties, particularly during the tenures of Presidents Biden and Trump. During his first term (2017–2021), Trump rolled back Obama-era fuel efficiency standards, slowing regulatory pressure on automakers to transition toward EVs. In his second term, the Trump administration has again taken a deregulatory approach. In contrast, the Biden administration adopted an aggressive pro-EV stance, prioritizing stricter fuel efficiency and emissions standards and allocating \$7.5 billion through the Infrastructure Investment and Jobs Act to develop a nationwide network of 500,000 EV charging stations.
@@
Table 3: Covariate effects (CE) from ordinal probit models. CE marked with $\times$ indicate parameters that are not statistically different from zero at 95% probability level.
@@
[…]
EV Info 0.0851-0.0851 0.0388-0.0388 $0.0217$ $0.1022$ .... .... .... .... .... .... .... ....
Env Better 0.1955-0.1955 0.0565-0.0565 $0.0970$ $0.1551$ .... .... .... .... .... .... .... ....
EV Infra .... .... .... .... .... .... .... .... 0.1880-0.1880 0.0553-0.0553 $0.0526$ $0.1908$
GACC (TM) $0.1971$ $0.0453$ 0.0945-0.0945 0.1478-0.1478 $0.3341$ $0.0133$ 0.1507-0.1507 0.1968-0.1968 $0.3170$ 0.0153-0.0153 0.1395-0.1395 0.1622-0.1622
GACC (RA) $0.1180$ $0.0391$ 0.0490-0.0490 0.1082-0.1082 $0.1649$ $0.0267$ 0.0639-0.0639 0.1278-0.1278 $0.1282$ $0.0079$ 0.0486-0.0486 0.0875-0.0875
EV Owner 0.1321-0.1321 0.0849-0.0849 $0.0051$ $0.2119$ 0.1749-0.1749 0.0794-0.0794 $0.0231$ $0.2312$ 0.1814-0.1814 0.0513-0.0513 $0.0421$ $0.1905$
Age <50<50 0.0326-0.0326 0.0130-0.0130 $0.0099$ $0.0356$ 0.0410-0.0410 0.0090-0.0090 $0.0129$ $0.0371$ 0.0371-0.0371 0.0036-0.0036 $0.0126$ $0.0282$
Female $\times$ $\times$ $\times$ $\times$ $0.0268$ $0.0055$ 0.0082-0.0082 0.0240-0.0240 $0.0301$ $0.0027$ 0.0100-0.0100 0.0228-0.0228
Grad & Above 0.0815-0.0815 0.0322-0.0322 $0.0276$ $0.0860$ 0.1139-0.1139 0.0240-0.0240 $0.0378$ $0.1001$ 0.1264-0.1264 0.0130-0.0130 $0.0445$ $0.0949$
Some College 0.0373-0.0373 0.0152-0.0152 $0.0099$ $0.0426$ 0.0541-0.0541 0.0118-0.0118 $0.0156$ $0.0503$ 0.0517-0.0517 0.0052-0.0052 0.0166-0.0166 $0.0403$
Married .... .... .... .... 0.0154-0.0154 0.0029-0.0029 $0.0050$ $0.0134$ $\times$ $\times$ $\times$ $\times$
Metropolitan 0.0534-0.0534 0.0179-0.0179 $0.0191$ $0.0522$ 0.0662-0.0662 0.0099-0.0099 $0.0232$ $0.0528$ 0.0512-0.0512 0.0027-0.0027 $0.0183$ $0.0356$
Midwest $0.0162$ $0.0061$ 0.0051-0.0051 0.0172-0.0172 $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
@@
West 0.0378-0.0378 0.0164-0.0164 $0.0101$ $0.0441$ 0.0457-0.0457 0.0109-0.0109 $0.0135$ $0.0431$ 0.0454-0.0454 0.0053-0.0053 $0.0148$ $0.0360$
US Born $0.0416$ $0.0186$ 0.0106-0.0106 0.0497-0.0497 $0.0777$ $0.0217$ 0.0214-0.0214 0.0781-0.0781 $0.0551$ $0.0071$ 0.0178-0.0178 0.0444-0.0444
Black $0.0614$ $0.0200$ 0.0220-0.0220 0.0594-0.0594 $0.0349$ $0.0058$ 0.0116-0.0116 0.0291-0.0291 $0.0460$ $0.0026$ 0.0160-0.0160 0.0326-0.0326
@@
Democrat 0.1211-0.1211 0.0578-0.0578 $0.0403$ $0.1386$ 0.1569-0.1569 0.0497-0.0497 $0.0513$ $0.1553$ 0.1882-0.1882 0.0353-0.0353 $0.0677$ $0.1558$
Ind & Others 0.0808-0.0808 0.0310-0.0310 $0.0219$ $0.0898$ 0.0890-0.0890 0.0187-0.0187 $0.0266$ $0.0811$ 0.1091-0.1091 0.0093-0.0093 $0.0356$ $0.0827$
[…]
@@
Income has a positive effect on EV purchase intent. Conditional on other factors, an additional \$50,000 in income changes the probability of strong interest (no interest) by 2.1–2.9 (2.2–3.6) percentage points. Our results find support in Zhao et al. (2022) and Mamkhezri (2025), who report greater support for EVs among higher-income individuals. The small covariate effect may reflect the growing availability of EVs across various market segments and the subsidies for low-income consumers, which reduce affordability barriers compared to earlier market years.
@@
Overall, the results show that interest in EVs varies by environmental attitudes, experience with EVs, confidence in charging infrastructure, and demographics. Some patterns likely reflect personal preferences; others may reflect barriers such as limited awareness or charging access.
@@ 4.2 Results from Ordinal Quantile Models
We now present the results from ordinal quantile models (Rahman, 2016) that shift attention from average probabilities of outcomes to those at specific quantiles, conditional on the covariates. To our knowledge, this aspect has not been examined in prior work on EVs. Table 4 presents the posterior means and standard deviations of the parameters from the ordinal quantile models estimated separately for each survey; along with the logarithm of marginal likelihood which can be used for Bayesian model comparison (Maheshwari and Rahman, 2023). Note that the coefficients in Table 4 are not directly comparable to those in Table 2 for two main reasons: (a) the models operate on different scales due to differences in the variance of the error terms, and (b) they address distinct inferential goals-the ordinal probit model focuses on the average probability of outcomes, whereas the ordinal quantile model targets the probability of outcomes at specific quantiles.
Table 5 reports the corresponding covariate effects for all variables whose parameters are statistically different from zero at 95% probability level. We see that covariate effects at the 50th percentile align closely with those from the ordinal probit model, but across all surveys, those at the 20th and 80th quantiles are noticeably different. This provides compelling evidence that the impact of covariates on purchase intent is not uniform but varies across the distribution of purchase intent, underscoring the value of quantile modeling in capturing such heterogeneity. Next, we emphasize the key differences across quantiles.
Although having heard or read extensively about EVs corresponds to higher purchase intent, the relationship is stronger among respondents more inclined toward adoption. In the 2021 survey and at the 80th percentile of purchase intent, respondents who have heard or read extensively about EVs are 12.6 percentage points more likely to express strong intent than those who have read or heard little to nothing. At the 50th and 20th percentiles of purchase intent, the difference is 9.9 and 6.1 percentage points, respectively.
Table 4: Results from quantile ordinal model: posterior mean (Mean) and standard deviation (Std) of the parameters, along with the logarithm of marginal likelihood ($\ln$ ML), from April 2021, May 2022, and May–Jun 2023 survey data.
@@
Intercept 3.60-3.60 $0.25$ 0.58-0.58 $0.17$ $2.60$ $0.25$ 2.15-2.15 $0.30$ $0.80$ $0.21$ $4.64$ $0.29$ 2.81-2.81 $0.30$ 0.05-0.05 $0.20$ $2.77$ $0.28$
EV Info $1.28$ $0.08$ $0.98$ $0.06$ $1.64$ $0.08$ .... .... .... .... .... .... .... .... .... .... .... ....
Env Better $2.52$ $0.11$ $1.80$ $0.07$ $2.85$ $0.10$ .... .... .... .... .... .... .... .... .... .... .... ....
EV Infra .... .... .... .... .... .... .... .... .... .... .... .... $2.83$ $0.13$ $1.92$ $0.10$ $2.71$ $0.14$
GACC (TM) 2.81-2.81 $0.13$ 1.96-1.96 $0.09$ 2.49-2.49 $0.13$ 3.79-3.79 $0.16$ 2.63-2.63 $0.10$ 3.70-3.70 $0.14$ 4.00-4.00 $0.17$ 2.41-2.41 $0.10$ 3.45-3.45 $0.14$
GACC (RA) 1.79-1.79 $0.10$ 1.24-1.24 $0.07$ 1.55-1.55 $0.11$ 1.94-1.94 $0.12$ 1.48-1.48 $0.08$ 2.03-2.03 $0.12$ 1.77-1.77 $0.12$ 1.15-1.15 $0.08$ 1.48-1.48 $0.12$
@@
Age<50<50 $0.40$ $0.08$ $0.36$ $0.05$ $0.58$ $0.08$ $0.50$ $0.09$ $0.37$ $0.07$ $0.63$ $0.10$ $0.42$ $0.11$ $0.34$ $0.07$ $0.55$ $0.10$
Female 0.11-0.11 $0.08$ 0.02-0.02 $0.05$ 0.03-0.03 $0.07$ 0.36-0.36 $0.10$ 0.22-0.22 $0.06$ 0.33-0.33 $0.09$ 0.49-0.49 $0.10$ 0.31-0.31 $0.07$ 0.27-0.27 $0.09$
Grad & Above $1.42$ $0.14$ 0.93-0.93 $0.08$ $1.17$ $0.13$ $1.93$ $0.16$ $1.04$ $0.10$ $1.19$ $0.13$ $1.94$ $0.16$ $1.09$ $0.10$ $1.43$ $0.13$
@@
Married .... .... .... .... .... .... $0.22$ $0.13$ $0.12$ $0.08$ $0.19$ $0.12$ $0.02$ $0.13$ $0.08$ $0.08$ $0.12$ $0.11$
@@
US Born 0.63-0.63 $0.12$ 0.49-0.49 $0.09$ 0.80-0.80 $0.14$ 0.89-0.89 $0.15$ 0.67-0.67 $0.10$ 1.25-1.25 $0.16$ 0.91-0.91 $0.16$ 0.45-0.45 $0.11$ 0.67-0.67 $0.15$
Black 1.23-1.23 $0.15$ 0.70-0.70 $0.10$ 0.56-0.56 $0.14$ 0.86-0.86 $0.20$ 0.25-0.25 $0.14$ 0.06-0.06 $0.18$ 0.94-0.94 $0.16$ 0.42-0.42 $0.11$ 0.36-0.36 $0.16$
Other Races $0.08$ $0.10$ $0.14$ $0.07$ $0.25$ $0.11$ 0.19-0.19 $0.14$ $0.07$ $0.10$ $0.22$ $0.13$ 0.30-0.30 $0.15$ $0.00$ $0.10$ $0.02$ $0.14$
@@
$\ln$ ML 12871-12871 12529-12529 12901-12901 8563-8563 8357-8357 8624-8624 8237-8237 7981-7981 8194-8194
Table 5: Covariate effects (CE) from quantile analysis. CE marked with $\times$ indicate parameters that are not statistically different from zero at 95% probability level.